# Efficient generative adversarial networks using linear additive-attention Transformers

### v1

@@ IV-D Lada discriminator
Using Transformers in the discriminator of GANs has been particularly challenging because regularization techniques such as R1 gradient penalty often lead to unstable training [7]. Although there are some notable exceptions to this (e.g. [7, 17]), they require laborious engineering and rarely outperform ConvNets. For this reason, most Transformer-based GANs employ ConvNet discriminators. We study the stability of Ladaformer discriminators trained with R1 gradient penalty compared to ConvNet discriminators, as well as the impact of the LEE and the bCR regularization.
Table II shows the FIDs obtained with different configurations of ConvNet and Ladaformer discriminators on CIFAR-10, CelebA and LSUN Bedroom. In addition, the number of FLOPs for the discriminator and generator, as well as the sizes of the embeddings in the generator corresponding to the three Ladaformer blocks in Figure 4, are also presented. As can be observed, the Ladaformer discriminators consistently outperform convolutional discriminators. However, the difference is small for CIFAR-10 and CelebA and significantly larger for LSUN Bedroom where larger generators are employed. Interestingly, a higher learning rate is required for the convolutional discriminator to match the Ladaformer discriminator FID. We also observe that, in LSUN Bedroom, using LEE leads to the lowest FID, while in CIFAR-10 and CelebA LEE does not make any difference. We hypothesize that this behavior is due to the resolution of the dataset, since high-resolution images require larger models. Note that in the LSUN Bedroom experiment, the Ladaformer blocks of the generator increase their embedding dimension locally from $\{1024,256,64\}$ to $\{1024,512,256\}$, which in turn increases the number of FLOPs
@@ IV-E Data efficiency
To analyze the behavior of LadaGAN under small-data regimes, we conducted experiments on CIFAR-10 with different training data sizes. In Table III, the FID scores of LadaGAN and StyleGAN2 with 10%, 20%, and 100% of training data are reported. As can be observed, LadaGAN outperforms StyleGAN2 in all scenarios without any hyperparameter changes. Moreover, LadaGAN with 20% of training data achieves a FID score relatively similar to StyleGAN2 with 100% of training data in both 50k and 10k evaluations. Note that since previous works compute the FID on CIFAR-10 using 10k sampled images (e.g. [46]) and we compute it using 50k sampled images, the results are not comparable. Therefore, we compute the FID scores using both 50k and 10k images and present them in Table III.
@@ IV-F Comparison with state-of-the-art models
We compare LadaGAN FID scores on CIFAR-10, CelebA, FFHQ, and LSUN Bedroom at two different resolutions with state-of-the-art single-step and multi-step image generation models. For a fair comparison, in addition to the evaluation described in subsection IV-B, we compute FID scores following VITGAN, SS-GAN, ADM-IP, and CT. More specifically, for CelebA, we compute the FID score between 19,962 generated samples and the 19,962 test images and also between 50k generated samples and the whole training set. For LSUN Bedroom $128\times 128$, we compute FID between 30k randomly sampled images and 30k generated images. For LSUN Bedroom $256\times 256$, we generate 50k images and use the same reference distribution statistics as CT, which was computed over 50k training samples. Consequently, we also apply the same data preprocessing for training such LadaGAN model.
@@
CT [33] 8.70$\ast$ - - - - 16.0
LadaGAN 3.48 2.91 1.81 4.48 5.08 6.36
Multi-step CT (2 steps) [33] 5.83$\ast$ - - - - 7.85
ADM-IP (80 steps) [6] 2.93 - 2.67 6.89 - -
ADM-IP (1000 steps) [6] 2.76 - 1.31 2.98 - -
On the other hand, we compare LadaGAN efficiency with state-of-the-art image generation models in terms of model size and complexity. Table V reports the number of parameters, FLOPs, throughput, and the number of images seen during training for LadaGAN, ViTGAN, StyleGAN2, CT, and ADM-IP with different image resolutions. For all datasets and resolutions, LadaGAN required the least number of parameters and FLOPs. In particular, LadaGAN required significantly fewer FLOPs: only 8.9% FLOPs of StyleGAN2 and 26.9% of ViTGAN for the $64\times 64$ resolution, and $\sim$37.5for the $128\times 128$ resolution. As expected, despite reducing the number of sampling steps, the number of FLOPs and throughput required for the multi-step generation models ADM-IP and CT is orders of magnitude higher than GANs, which are single-step generators. Note that although for the $32\times 32$ resolution ADM-IP and LadaGAN require almost the same number of images during training, when the resolution increases ($64\times 64$) LadaGAN requires approximately half the images of ADM-IP. Finally, we find that CT has significantly more parameters than LadaGAN and is the model that requires to train longer.
Remarkably, in contrast to ADM-IP which drastically increases the number of parameters and FLOPs for higher resolutions, LadaGAN parameters and FLOPs remain practically the same between $32\times 32$ and $64\times 64$; this is because instead of generating pixels, the final Ladaformer block generates patches of $2\times 2$, resulting in almost the same architecture.
IV-G Generated images and interpolation
@@
Figure 6: CIFAR-10 (a), CelebA (b), LSUN Bedroom (c), and FFHQ (d) with corresponding additive attention maps for a single head of the $32\times 32$, $16\times 16$, and $8\times 8$ Ladaformer blocks.
@@ V Conclusion
In this paper, we presented LadaGAN, an efficient GAN architecture based on a novel linear additive-attention block called Ladaformer. This block showed to be more suitable for both the generator and the discriminator than other efficient Transformer blocks, allowing stable GAN training in different scenarios. Our findings indicate that Ladaformer is convolution-compatible and LadaGAN is gradient-stable and highly effective for image generation tasks. Remarkably, LadaGAN outperformed ConvNet and Transformer GANs on multiple benchmark datasets at different resolutions while requiring significantly fewer FLOPs. Moreover, compared with diffusion models and Consistency Training, LadaGAN achieves competitive performance at a fraction of the computational cost.
@@
As future work, we plan to train LadaGAN in audio and text-to-image scenarios. Moreover, the Ladaformer block and its compatibility with convolutions remain to be explored in other tasks, like image and video classification.

### v4

@@ IV-D Lada discriminator
Using Transformers in the discriminator of GANs has been particularly challenging because regularization techniques such as $R_{1}$ gradient penalty often lead to unstable training [7]. Although there are some notable exceptions to this (e.g. [7, 17]), they require laborious engineering and rarely outperform ConvNets. For this reason, most Transformer-based GANs employ ConvNet discriminators. In this context, we study the stability of Ladaformer discriminators trained with $R_{1}$ gradient penalty compared to ConvNet discriminators, as well as the impact of increasing the model size with LEE.
Table II shows the FIDs obtained with different configurations of ConvNet and Ladaformer discriminators on CIFAR-10, CelebA and LSUN Bedroom. In addition, the number of FLOPs for the discriminator and generator, as well as the sizes of the embeddings in the generator corresponding to the three Ladaformer blocks in Figure 4, are also presented. As can be observed, the Ladaformer discriminators consistently outperform convolutional discriminators. However, the difference is small for CIFAR-10 and CelebA and significantly larger for LSUN Bedroom where larger generators are employed. Interestingly, a higher learning rate is required for the convolutional discriminator to match the Ladaformer discriminator FID. We also observe that, in LSUN Bedroom, using LEE leads to the lowest FID, while in CIFAR-10 and CelebA LEE does not make any difference. We hypothesize that this behavior is due to the resolution of the dataset, since high-resolution images require larger models. Note that in the LSUN Bedroom experiment, the Ladaformer blocks of the generator increase their embedding dimension locally from $\{1024,256,64\}$ to $\{1024,512,256\}$, which in turn increases the number of FLOPs. deteriorates
@@ IV-E Data efficiency
To analyze the behavior of LadaGAN under small-data regimes, we conducted experiments on CIFAR-10 with different training data sizes with and without bCR regularization. In Table III, the FID scores of LadaGAN and StyleGAN2 with 10%, 20%, and 100% of training data are reported. As can be observed, LadaGAN with bCR outperforms StyleGAN2 in all scenarios without any hyperparameter changes. Moreover, LadaGAN with 20% of training data achieves a FID score relatively similar to StyleGAN2 with 100% of the training data in both 50k and 10k evaluations. On the other hand, without bCR LadaGAN performance declines in the 10% and 20% data regimes, where is outperformed by StyleGAN2. However, for the complete data regime the performance drop is less pronounced, outperforming StyleGAN2. These results show that LadaGAN benefits from bCR when a small training dataset is available, although these benefits decrease as the dataset size increases. Note that since previous works compute the FID on CIFAR-10 using 10k sampled images (e.g. [46]) and we compute it using 50k sampled images, the results are not comparable. Therefore, we compute the FID scores using both 50k and 10k images and present them in Table III.
@@ IV-F Comparison with state-of-the-art models
We compare LadaGAN FID scores on CIFAR-10, CelebA, FFHQ, and LSUN Bedroom at two different resolutions with state-of-the-art single-step and multi-step image generation models. For a fair comparison, in addition to the evaluation described in subsection IV-B, we compute FID scores following VITGAN, SS-GAN, ADM-IP, and CT. More specifically, for CelebA, we compute the FID score between 19,962 generated samples and the 19,962 test images and also between 50k generated samples and the whole training set. For LSUN Bedroom $128\times 128$, we compute FID between 30k randomly sampled images and 30k generated images. For LSUN Bedroom $256\times 256$, we generate 50k images and use the same reference distribution statistics as CT, which is computed over 50k training samples. Consequently, we also apply the same data preprocessing for training such LadaGAN model.
@@
CT [33] 8.70$\ast$ - - - - 16.0$\ast$
LadaGAN 3.29 2.89 1.81 4.48 5.08 6.36
Multi-step CT (2 steps) [33] 5.83$\ast$ - - - - 7.85$\ast$
ADM-IP (80 steps) [6] 2.93$\ast$ - 2.67$\ast$ 6.89$\ast$ - -
ADM-IP (1000 steps) [6] 2.76$\ast$ - 1.31$\ast$ 2.98$\ast$ - -
@@
IV-G Computational cost analysis
We compare LadaGAN efficiency with state-of-the-art image generation models in terms of model size and complexity. Table V reports the number of parameters, FLOPs, throughput, and the number of images seen during training for LadaGAN, ViTGAN, StyleGAN2, CT, and ADM-IP with different image resolutions. For all datasets and resolutions, LadaGAN required the least number of parameters and FLOPs. In particular, LadaGAN required significantly fewer FLOPs: only 8.9% FLOPs of StyleGAN2 and 26.9% of ViTGAN for the $64\times 64$ resolution, and $\sim$37.5for the $128\times 128$ resolution. As expected, despite reducing the number of sampling steps, the number of FLOPs and throughput required for the multi-step generation models ADM-IP and CT is orders of magnitude higher than GANs, which are single-step generators. Note that although for the $32\times 32$ resolution ADM-IP and LadaGAN require almost the same number of images during training, when the resolution increases ($64\times 64$) LadaGAN requires approximately half the images of ADM-IP. Finally, we find that CT has significantly more parameters than LadaGAN and is the model that requires to train longer.
Remarkably, in contrast to the ADM-IP CelebA ($64\times 64$) model, which requires 5 days for training on 16 Tesla V100 GPUs (16G memory for each GPU) [6], LadaGAN requires only 35 hours using a single RTX 3080 Ti GPU (12G memory) to observe the same number of images. Even for CIFAR 10 ($32\times 32$), ADM-IP takes 2 days using 2 GPUs, whereas LadaGAN is trained in less than 35 hours on a single GPU.
Finally, LadaGAN parameters and FLOPs remain practically the same between $32\times 32$ and $64\times 64$; this is because instead of generating pixels, the final Ladaformer block generates patches of $2\times 2$, resulting in almost the same architecture.
IV-H Generated images and interpolation
Figure 6: Samples from LadaGAN models on CIFAR-10 (a), CelebA (b), LSUN Bedroom (c), and FFHQ (d), along with corresponding additive attention maps for a single head from the $32\times 32$, $16\times 16$, and $8\times 8$ Ladaformer blocks.
@@ V Conclusion
In this paper, we presented LadaGAN, a novel efficient GAN architecture based on a linear additive-attention block called Ladaformer. This block showed to be more suitable for both the generator and the discriminator than other efficient Transformer blocks, allowing stable GAN training in different scenarios. Our findings indicate that Ladaformer is convolution-compatible and LadaGAN is gradient-stable and highly effective for image generation tasks. Remarkably, LadaGAN outperformed ConvNet and Transformer GANs on multiple benchmark datasets at different resolutions while requiring significantly fewer FLOPs. Moreover, compared with diffusion models and Consistency Training, LadaGAN achieves competitive performance at a fraction of the computational cost.
@@
As future work, we plan to train LadaGAN in audio and text-to-image scenarios, and more diverse datasets. Moreover, the difference in efficiency and FID evaluation between using patch generation or convolutional decoders when increasing the image resolution and dataset size remains to be studied. Finally, we believe that the Ladaformer block and its compatibility with convolutions are worth exploring in other tasks, like image and video classification.